OpenStudy (anonymous):

i need an example for my homework. please show me every step. 4t-7-t+=7-2t-5

OpenStudy (anonymous):

its algebra you can combine same variables you can do operations to one side so long as you do them to the other

OpenStudy (anonymous):

combine the t's then get all the t's on one side and all the numbers on the other

OpenStudy (anonymous):

i understand that but i am having trouble wrapping my head around this problem.

OpenStudy (anonymous):

4t-7-t+=7-2t-5 3t-7=2-2t 5t-7=2 5t=9 t=9/5

OpenStudy (mertsj):